OpenStudy (anonymous):

Determine the value of k such that g(x) = 3x + k intersects the quadratic function f(x) = 2x^2 - 5x + 3

OpenStudy (anonymous):

The answer is supposed to be \[k \ge -5\]

OpenStudy (anonymous):

and after I equated I got 2x^2 - 8x + 3-k

sam (.sam.):

b^2 -4ac=0

OpenStudy (anonymous):

and then I use the discriminat: (-8)^2 - 4(2)(3-k) = 64 - 8 - 12 but I don't know how to get the k.. is it +4k?

sam (.sam.):

a=2 b=-8 c=3-k

OpenStudy (hoblos):

f(x) = g(x) 2x^2 - 5x + 3 = 3x + k 2x^2 -8x +3-k = 0 the two graphs intersect when the quadratic has real solutions 64 - 8(3-k) >= 0 40 +8k >=0 k >= -5

sam (.sam.):

(-8)^(2)-4(2)(3-k)=0 ((64))-4(2)(3-k)=0 (64)-4(2)(3-k)=0 64-4(2)(3-k)=0 64-4(6-2k)=0 64-24+8k=0 40+8k=0 8k=-40 k=-5

OpenStudy (anonymous):

ohhh so I multiply the ones with the brackets first?

sam (.sam.):

yes